# Does provable absence of barren plateaus imply classical simulability? Or, why we need to rethink variational quantum computing

### v2

@@ Appendix B Examples of non-concentrated but also non-simulable loss functions
for all $x\in\{0,1\}^{n}$, where $Z_{1}$ is the Pauli-Z operator acting on the first qubit. As previously stated, we will focus on the fact that $B(x)$ is classically hard to simulate for uniformly random input bitstrings $x\in\{0,1\}^{n}$, even given a polynomial-size advice string. Equivalently, this can be stated as saying that this class of Boolean functions $\{B(x)\}$ is not in $\mathrm{BPP}/\mathrm{poly}$. As data from quantum experiments are a restricted form of advice strings, we will use this observation to argue that a loss function that estimates $B(x)$ for $x$ uniformly sampled from $\{0,1\}^{n}$ is not ${\rm CSIM}_{\mathrm{QE}}$.
@@
$\ell_{\bm{\theta}}(O)=\bra{0^{n}}\left(\bigotimes_{i=1}^{n}e^{i\theta_{i}X}\right)U^{\dagger}Z_{1}U\left(\bigotimes_{i=1}^{n}e^{-i\theta_{i}X}\right)\ket{0^{n}}\,,$ (52)
where $O=Z_{1}$ and we have explicitly taken the initial state $\rho$ to be the all-zero state. Let us further suppose that each $\theta_{i}$ is uniformly sampled from the discrete set $\{0,\pi/2,\pi,3\pi/2,2\pi\}$ such that the parametrized quantum circuit prepares a random computational basis state. That is, $\left(\bigotimes_{i=1}^{n}e^{-i\theta_{i}X}\right)\ket{0^{n}}=\ket{x}$ where $|x\rangle$ is a random bitstring. We call this distribution $\mathcal{P}_{\rm disc}$. We will then define $\mathcal{C}_{{\rm Bool}}^{\rm disc.}$ as the class of all instances where the circuit is of the form in Eq. (50) where $U$ implements a classically hard Boolean function (even with advice strings) and $\bm{\theta}\sim\mathcal{P}_{\rm disc}$.
Since this circuit by assumption approximates $B(x)$ for all $x$, as stated in Eq. (51), and estimating $\{B(x)\}$ for uniformly sampled $x$ is not in $\mathrm{BPP}/\mathrm{poly}$, and data from quantum experiments are a form of advice string, it follows that $\mathcal{C}_{{\rm Bool}}^{\rm disc.}$ is not in ${\rm CSIM}_{\mathrm{QE}}$. Moreover, since $U$ is a polynomial-size circuit, the problem is in ${\rm QSIM}$. That is, we have
@@
where $\overline{{{\rm CSIM}_{\mathrm{QE}}}}$ denotes the classes not in ${\rm CSIM}_{\mathrm{QE}}$.
@@
Thus, $\mathcal{C}_{{\rm Bool}}^{\rm disc.}$ provides an example of a problem class that is in $\overline{{\rm BP}}$ but not in ${\rm CSIM}_{\mathrm{QE}}$, as claimed.
However, this example is rather contrived. For one, it considers an initialization over a discrete set of parameters. This is a long way from the spirit of more variational quantum algorithms. Moreover, this discrete initialization is doing much of the heavy lifting in this example. To see this one can also consider a variant of the above problem where the parameters are sampled uniformly in the range $[0,2\pi]$. We will call this distribution $\mathcal{P}_{\rm cont.}$ and the corresponding problem class $\mathcal{C}_{{\rm Bool}}^{\rm cont.}$.
@@
It then follows from Claim 2 that $\mathcal{C}_{{\rm Bool}}^{\rm cont.}$ is classically simulable in the sense that with high probability for $\bm{\theta}\sim\mathcal{P}_{\rm cont.}$ one can provide a classical estimate of the loss defined in Eq. (52).
@@
So what do we conclude from these examples? Firstly, this is a concrete example of the limitations of an average case notion of simulability. The loss we have constructed here is easy to classically simulate on average over the continuous parameter range $\mathcal{P}_{\rm cont.}$ after an initial data acquisition phase. However, on a special measure-zero set of parameters, the problem becomes classically hard. Secondly, while in a typical random product state the expectation of any given global operators vanishes with high probability, something subtle can happen for the case of random computational basis states. Namely, while on average the contribution of any given global Pauli is exponentially small, there are exponentially many such terms and with the right structure these can add up to give a non-zero total contribution. This is what happens for $\mathcal{C}_{{\rm Bool}}^{\rm dist.}$- the contribution from $A^{H}$ does not vanish and gives rise to a non-vanishing variance. However, in parallel, the exponential number of global terms in $A^{H}$ makes its contribution impossible to classically simulate. This suggests smart initialization procedures may be the key to constructing efficiently trainable variational quantum algorithms that are not also classically simulable.

### v3

@@ A.8 Parameterized Dynamic Circuit
for all $x\in\{0,1\}^{n}$, where $Z_{1}$ is the Pauli-Z operator acting on the first qubit. As previously stated, we will focus on the fact that $B(x)$ is classically hard to simulate for uniformly random input bitstrings $x\in\{0,1\}^{n}$, even given a polynomial-size advice string. Equivalently, this can be stated as saying that this class of Boolean functions $\{B(x)\}$ is not in $\mathrm{BPP}/\mathrm{poly}$. As data from quantum experiments are a restricted form of advice strings, we will use this observation to argue that a loss function that estimates $B(x)$ for $x$ uniformly sampled from $\{0,1\}^{n}$ is not $\rm QESIM$}.
@@
[…]
Since this circuit by assumption approximates $B(x)$ for all $x$, as stated in Eq. (65), and estimating $\{B(x)\}$ for uniformly sampled $x$ is not in $\mathrm{BPP}/\mathrm{poly}$, and data from quantum experiments are a form of advice string, it follows that $\mathcal{C}_{{\rm Bool}}^{\rm disc.}$ is not in $\rm QESIM$}. Moreover, since $U$ is a polynomial-size circuit, the problem is in $\rm QSIM$}. That is, we have
@@
where $\overline{{{\rm CSIM}_{\mathrm{QE}}}}$ denotes the classes not in $\rm QESIM$}.
@@
Thus, $\mathcal{C}_{{\rm Bool}}^{\rm disc.}$ provides an example of a problem class that is in $\overline{{\rm BP}}$ but not in $\rm QESIM$}, as claimed.
However, this example is rather contrived. For one, it considers an initialization over a discrete set of parameters. This is a long way from the spirit of more variational quantum algorithms. Moreover, this discrete initialization is doing much of the heavy lifting in this example. To see this one can also consider a variant of the above problem where the parameters are sampled uniformly in the range $0,2\pi]$. We will call this distribution $\mathcal{P}_{\rm cont.}$ and the corresponding problem class $\mathcal{C}_{{\rm Bool}}^{\rm cont.}$.
@@
It then follows from Claim 2 in the main text that $\mathcal{C}_{{\rm Bool}}^{\rm cont.}$ is classically simulable in the sense that with high probability for $\bm{\theta}\sim\mathcal{P}_{\rm cont.}$ one can provide a classical estimate of the loss defined in Eq. (66).
@@
So what do we conclude from these examples? Firstly, this is a concrete example of the limitations of an average case notion of simulability. The loss we have constructed here is easy to classically simulate on average over the continuous parameter range $\mathcal{P}_{\rm cont.}$ after an initial data acquisition phase. However, on a special measure-zero set of parameters, the problem becomes classically hard. Secondly, while in a typical random product state the expectation of any given global operator vanishes with high probability, something subtle can happen for the case of random computational basis states. Namely, while on average the contribution of a typical global Pauli is exponentially small, there are exponentially many such terms and with the right structure these can add up to give a non-zero total contribution. This is what happens for $\mathcal{C}_{{\rm Bool}}^{\rm dist.}$—the contribution from $A^{H}$ does not vanish and gives rise to a non-vanishing variance. However, in parallel, the exponential number of global terms in $A^{H}$ makes its contribution impossible to classically simulate. This suggests smart initialization procedures may be the key to constructing efficiently trainable variational quantum algorithms that are not also classically simulable.
Supp. Info. C Absence of barren plateaus in shallow local circuits with arbitrary topology
In this section we will show that that shallow hardware efficient ansatz composed of two-qubit gates acting an arbitrary topology do not posses barren plateaus. While these results can be, in a way, derived from those in [106, 107], we add the full proof here for completeness.
We begin by present a few useful definitions. First, let $\mathcal{H}=(\mathbb{C}^{2})^{\otimes 2}$ be the Hilbert space of two qubits. In what follows we will denote the subsystem of the first qubit with an index $A$, and that of the second qubit with an index $B$ so that $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Then let $\mathcal{B}(\mathcal{H}^{\otimes t})$ be the set of bounded operators acting on $t$ copies of the aforementioned Hilbert space. We are interested in computing expectation values of the form
$\tau^{(t)}_{U}(X)=\int_{SU(4)}d\mu(U)\,U^{\otimes t}X(U^{\dagger})^{\otimes t}\,$, (71)
where $d\mu(U)$ denotes the uniquely defined, normalized, Haar measure over the fundamental representation of $SU(4)$, and where $X\in\mathcal{B}(\mathcal{H}^{\otimes t})$.
[…]
$\mathcal{H}^{\otimes 2}=(\mathcal{H}_{A,1}\otimes\mathcal{H}_{B,1})\otimes(\mathcal{H}_{A,2}\otimes\mathcal{H}_{B,2})\rightarrow(\mathcal{H}_{A,1}\otimes\mathcal{H}_{A,2})\otimes(\mathcal{H}_{B,1}\otimes\mathcal{H}_{B,2})\,$, (75)
so that
[…]
where we defined
[…]
Combining the previous results, we obtain
[…]
From here, we can find that an orthonormal basis for the commutant of the two-fold tensor rep of $SU(4)$ is
[…]
where we have dropped the trivial identity terms for ease of notation. Then, we can write the two-fold twirl as
[…]
Given Eq. (80) we can now compute the following twirls:
$\tau^{2}_{U}(P_{A}\otimes P_{B})$ $=\frac{1}{15}({\rm SW}_{A}+{\rm SW}_{B}+{\rm SW}_{A}\otimes{\rm SW}_{B})\,$, (81)
$\tau^{2}_{U}({\rm SW}_{A})$ $=\frac{1}{5}({\rm SW}_{A}+{\rm SW}_{B}+{\rm SW}_{A}\otimes{\rm SW}_{B})\,$, (82)
$\tau^{2}_{U}({\rm SW}_{B})$ $=\frac{1}{5}({\rm SW}_{A}+{\rm SW}_{B}+{\rm SW}_{A}\otimes{\rm SW}_{B})\,$, (83)
$\tau^{2}_{U}({\rm SW}_{A}\otimes{\rm SW}_{B})$ $=\frac{3}{5}({\rm SW}_{A}+{\rm SW}_{B}+{\rm SW}_{A}\otimes{\rm SW}_{B})\,$, (84)
where $P_{A(B)}$ is a single-qubit Pauli acting on the subsystems $A(B)$.
Supplementary Figure 6: Four qubit circuit example. Consider the following circuit acting on four qubits. At the end of the circuit we measure a Pauli operator on qubits 2 and 3.
[…]
$\mathbb{E}_{U}[f(U)]$ $=\int_{SU(4)}d\mu(U_{12})d\mu(U_{34})d\mu(U_{23})f(U)\,$, (85)
Using Eq. (72), we can readily show that
[…]
Then, since
$\Tr[U\rho U^{\dagger}O]^{2} =Tr$$=\Tr[U^{\otimes 2}\rho^{\otimes 2}(U^{\dagger})^{\otimes 2}O]$
=Tr[$=\Tr[\rho^{\otimes 2}(U^{\dagger})^{\otimes 2}OU^{\otimes 2}]\,$,
we find
$\mathbb{E}_{U}[\Tr[\rho^{\otimes 2}(U^{\dagger})^{\otimes 2}OU^{\otimes 2}]] =Tr$$=\Tr[\rho^{\otimes 2}\mathbb{E}_{U}[(U^{\dagger})^{\otimes 2}OU^{\otimes 2}]]\,$,
where we have used the linearity of the trace. Then, we see that we need to compute
$\tau_{U_{12}}^{(2)}\circ\tau_{U_{34}}^{(2)}\circ\tau_{U_{23}}^{(2)}[O^{\otimes 2}]\,.$ (87)
Since $O$ is some Pauli acting non trivially on qubits $2$ and $3$, we find via Eq. (81)
$\tau_{U_{23}}^{(2)}[O^{\otimes 2}]=\frac{1}{15}({\rm SW}_{2}+{\rm SW}_{3}+{\rm SW}_{2}\otimes{\rm SW}_{3})\,.$ (88)
We then use Eq. (82)–(84) to obtain
$\tau_{U_{34}}^{(2)}[\frac{1}{15}({\rm SW}_{2}+{\rm SW}_{3}+{\rm SW}_{2}\otimes{\rm SW}_{3})]$
$=$ $\frac{{\rm SW}_{2}}{15}+\frac{1}{75}({\rm SW}_{3}+{\rm SW}_{4}+{\rm SW}_{3}\otimes{\rm SW}_{4})+\frac{3}{75}({\rm SW}_{2}\otimes{\rm SW}_{3}+{\rm SW}_{2}\otimes{\rm SW}_{4}+{\rm SW}_{2}\otimes{\rm SW}_{3}\otimes{\rm SW}_{4})\,.$ (89)
Here, we can already observe an important phenomenon. If a ${\rm SW}_{i}$ goes into a twirl from a gate $U_{ij}$, then we obtain terms ${\rm SW}_{i}$, ${\rm SW}_{j}$, and ${\rm SW}_{i}\otimes{\rm SW}_{j}$ with a coefficient $\frac{1}{5}$. Similarly, if ${\rm SW}_{i}\otimes{\rm SW}_{j}$ goes into a twirl from a gate $U_{ij}$, then we obtain terms ${\rm SW}_{i}$, ${\rm SW}_{j}$, and ${\rm SW}_{i}\otimes{\rm SW}_{j}$ with a coefficient $\frac{3}{5}$. Hence, we simply need to keep track of indexes and coefficients. As such, let us we-write Eq. (88) as
$O^{\otimes 2}\xrightarrow[\tau_{U_{23}}^{(2)}]{}\left\{\left\{(2),\frac{1}{5}\right\},\left\{(3),\frac{1}{5}\right\},\left\{(2,3),\frac{1}{5}\right\}\right\}\,.$ (90)
Then, the next twirl over $U_{34}$ in Eq. (89) leads to
$\xrightarrow[\tau_{U_{33}}^{(2)}$]{} $\Big{\{}\left\{(2),\frac{1}{5}\right\},\left\{(3),\frac{1}{75}\right\},\left\{(4),\frac{1}{75}\right\},\left\{(3,4),\frac{1}{75}\right\},$
$\left\{(2,3),\frac{3}{75}\right\},\left\{(2,4),\frac{3}{75}\right\},\left\{(2,3,4),\frac{3}{75}\right\}\Big{\}}$ (91)
and finally the next twirl over $U_{12}$
$\xrightarrow[\tau_{U_{33}}^{(2)}$]{} $\Big{\{}\left\{(1),\frac{1}{75}\right\},\left\{(2),\frac{1}{75}\right\},\left\{(3),\frac{1}{75}\right\},\left\{(4),\frac{1}{75}\right\},$
$\left\{(1,2),\frac{1}{75}\right\},\left\{(3,4),\frac{1}{75}\right\},\left\{(2,3),\frac{3}{375}\right\},$
$\left\{(1,3),\frac{3}{375}\right\},\left\{(2,4),\frac{3}{375}\right\},\left\{(1,4),\frac{3}{375}\right\},$
$\left\{(1,2,3),\frac{3}{375}\right\},\left\{(1,2,4),\frac{3}{375}\right\},\left\{(2,3,4),\frac{3}{375}\right\},$
$\left\{(1,3,4),\frac{3}{375}\right\},\left\{(1,2,3,4),\frac{3}{375}\right\}\Big{\}}\,.$ (92)
The previous example unveils a simple pattern that will be universally true. First, the “most-local” operators, are the ones which will be suppressed the least. Whereas “more global” operators will be quickly suppressed. As such, we can intuitively see that, in average, the Heisenberg evolved operator will live on $\mathcal{O}(1)$ bodied operators. Indeed, one can study how much the coefficient of a given operators at the out put of the circuit will look like by evaluating some weighted random walk over the circuit architecture following the rules mentioned above.
Putting it all together, we are now ready to show that shallow local circuits avoid barren plateaus when measured locally at their output. Interestingly, this result encompassed, and constitutes a simpler proof, to those in Refs. [24] and [15]. Consider therefore an arbitrary circuit with two-qubit gate topology
$\mathcal{A}=\{(i_{k},j_{k}),\quad\forall k=1,\ldots,K\}\,.$ (93)
Note that we are not assuming any dimensionality on $\mathcal{A}$, meaning that it could be as general as desired. Then, if the measurement operator is locally acting on the $\mu$-th qubit, and that this index appears in $\mathcal{A}$ $L$ times, one can lower bound the expectation value’s variance as
[…]
Notably, we find that the reduced states will be exponentially close to the maximally mixed state if they satisfy a volume law of entanglement, whereas their distance to it will be polynomially small is they follow an area law of entanglement [42].